OpenStudy (anonymous):

Which of the following is equivalent to 16a2 − 49? a. (4a − 7) (4a − 7) b. (4a + 7) (4a − 7) c. (7 − 4a) (7 − 4a) d. (7 + 4a) (7 − 4a)

OpenStudy (anonymous):

\[x^2-y^2=(x+y)(x-y)\] in this case \(x=4a, y = 7\)

OpenStudy (anonymous):

so the answer is c

OpenStudy (anonymous):

no

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so what is it?

OpenStudy (anonymous):

What @satellite73 gave you, he gave you the answer, base on that. Try to connect his answer to your problem.

OpenStudy (anonymous):

so b?

OpenStudy (anonymous):

Correct.

OpenStudy (anonymous):

thanks